I don't understand this problem. Consider a gameboard that has 5x5=25 squares. You ou one grain of sand on the first square, two grains on the second square, 4 grains on the third square, continuing until you have filled the board. Compute the total number of grains needed to do this to see why the task, seemingly simple, can not be completed. The three multiple choice answers were: 1.68x10^7 grains of sand, 3.36x10^7 grains of sand, or 1.00x10^6 grains of sand?
Answer by stanbon(75887)

Consider a gameboard that has 5x5=25 squares. You ou one grain of sand on the first square, two grains on the second square, 4 grains on the third square, continuing until you have filled the board. Compute the total number of grains needed to do this to see why the task, seemingly simple, can not be completed. The three multiple choice answers were: 1.68x10^7 grains of sand, 3.36x10^7 grains of sand, or 1.00x10^6 grains of sand?
-----
Sun = 2^0+2^1*2^2+...+ 2^24 = 2^0(2^25 -1)/(2-1) = 2^25-1 = 3.36x10^7
==================
Cheers,
Stan H.